OpenStudy (anonymous):

WILL FAN AND MEDAL!! In complete sentences, explain why and how to simplify the following radicals: 1) √63 2) √96 3) √x^23 4) √40x^5y

ganeshie8 (ganeshie8):

heard of `prime factorization` before ?

ganeshie8 (ganeshie8):

\(\large \sqrt{63} = \sqrt{9 \times 7} = \sqrt{3^2\times 7} = 3\sqrt{7}\)

ganeshie8 (ganeshie8):

Step1 : The goal here is to find the prime factors of 63. Since 63 is not prime, find two numbers that multiply to 63. How about 9 and 7? Since 7 is prime, it doesn’t need to be factored anymore. However, 9 may be written as 3 times 3. Three is also a prime number. Therefore, the prime factors of 63 are 7, 3, and 3.

ganeshie8 (ganeshie8):

Step2 : Next you need to look for pairs of factors. For 63, there is only one pair of factors inside the radical: 3 times 3. Since the square root of 3 times 3 is the same as 3, a pair of factors inside the radical becomes a single factor outside the radical. Any remaining factors stay inside the radical. The square root of 63 simplifies to 3 times the square root of 7.

ganeshie8 (ganeshie8):

See if you can work the remaining problems :)